# Exploring the Most Extreme Gamma-Ray Blazars Using Broadband Spectral Energy Distributions

### v1

@@ 6 Summary and conclusions
The broadband SEDs were modelled assuming non-flaring states, allowing us to determine key parameters such as synchrotron peak frequency ($\nu_{sync}^{peak}$) and CD. Using a one-zone SSC model, complemented by host galaxy templates, we identified 66 new EHSP candidates, approximately doubling the known population of these rare and intriguing objects. The results revealed that most EHSPs in the sample present low CD values (CD $<1$), consistent with SSC-dominated emission in environments with weak external photon fields. Host galaxy emission was fitted using templates for lenticular and elliptical galaxies, but no significant differences were observed in the radiative properties of sources hosted by different galaxy types. This suggests that the thermal contribution from the host galaxy has a negligible impact on the high-energy emission.
@@ Acknowledgements.
M.L. and J.L.C acknowledge the support of MCIN project PID2022-138132NB-C42 and grant PRE2020-093502. M.N.R. acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant PARTICIPACIÓN DEL IAC EN EL EXPERIMENTO AMS and the European Regional Development Fund (ERDF) with reference PID2022- 137810NB-C22/DIO 10.13039/501100011033. A.D. is thankful for the support of Proyecto PID2021-126536OA-I00 funded by MCIN / AEI /10.13039/501100011033. This research has made use of the CTA instrument response functions provided by the CTA Consortium and Observatory, see https://www.ctao-observatory.org/science/cta-performance/ (version prod5 v0.1; Cherenkov Telescope Array Observatory and Cherenkov Telescope Array Consortium, 2021) for more details. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. Part of this work is based on archival data, software or online services provided by the Space Science Data Center - ASI.

### v2

@@ 6 Summary and conclusions
The broadband SEDs were modelled assuming non-flaring states, allowing us to determine key parameters such as synchrotron peak frequency ($\nu_{sync}^{peak}$) and CD. Using a one-zone SSC model, complemented by host galaxy templates, we identified 66 new EHSP candidates, approximately doubling the known population of these rare and intriguing objects. The results revealed that most EHSPs in the sample present low CD values (CD <1<1< 1), consistent with SSC-dominated emission in environments with weak external photon fields. Host galaxy emission was fitted using templates for lenticular and elliptical galaxies, but no significant differences were observed in the radiative properties of sources hosted by different galaxy types. This suggests that the thermal contribution from the host galaxy has a negligible impact on the high-energy emission.
@@
Data availability
Tables 2, LABEL:table:bestfit and LABEL:table:energy-budget are available at the CDS via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
M.L. and J.L.C acknowledge the support of MCIN project PID2022-138172NB-C42 and grant PRE2020-093502. M.N.R. acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant PARTICIPACIÓN DEL IAC EN EL EXPERIMENTO AMS and the European Regional Development Fund (ERDF) with reference PID2022- 137810NB-C22/DIO 10.13039/501100011033. A.D. is thankful for the support of Proyecto PID2021-126536OA-I00 funded by MCIN / AEI /10.13039/501100011033. This research has made use of the CTA instrument response functions provided by the CTA Consortium and Observatory, see https://www.ctao-observatory.org/science/cta-performance/ (version prod5 v0.1; Cherenkov Telescope Array Observatory and Cherenkov Telescope Array Consortium, 2021) for more details. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. Part of this work is based on archival data, software or online services provided by the Space Science Data Center - ASI.
@@ Appendix B Best-fit parameters and energetic report of the sample sources
Appendix C FITS File Structure
This section summarizes the contents of the FITS file published in Zenodo (doi: 10.5281/zenodo.15778399) and in https://www.ucm.es/blazars/ehsp. The HDUs contained in the FITS file are summarized in Table 5. Table 1 contains the best-fit parameter values obtained from the broadband SED modelling performed for the 124 sources in the sample. Table 2 contains the resulting synchrotron peak and IC peak frequencies, Compton dominance, jet luminosities of each component, and the magnetic to kinetic energy density ratio for the 124 sources in the sample. Table 3 includes the expected detection significance with CTAO for all the sources, as well as their redshift. Table 4 contains for each source in the sample all the multi-wavelength data that has been used to reconstruct the broadband SED and to perform the SED modelling. The description of the contents included in these tables is given in Tables 6, 7, 8 and 9.
HDU Index Type Shape Description
1 BinTableHDU 124R x 15C Best-fit parameters resulting from the SED modelling
2 BinTableHDU 124R x 10C Other derived physical parameters and energetic report
3 BinTableHDU 124R x 4C CTAO expected detection significance
4 BinTableHDU 124R x 6C Multi-wavelength SED data used for the modelling
Table 5: HDUs contained in the FITS file.
Column Name Data Type Units Description
‘Source_Name’ bytes16 - Source name given in the 4FGL catalogue
‘B’ float64 G Magnetic field strength
‘R’ float64 cm Radius of the emitting region
‘R_H’ float64 cm Distance from the emitting region to the central black hole
‘theta’ float64 deg Jet viewing angle
‘BulkFactor’ float64 - Bulk Lorentz factor of the electrons in the jet
‘gmin’ float64 - Minimum Lorentz factor of the electron population
‘gmax’ float64 - Maximum Lorentz factor of the electron population
‘N’ float64 1/cm3 Electron density
‘gamma_break’ float64 - Break Lorentz factor
‘p1’ float64 - Spectral slope of the lower energy electron population
‘p2’ float64 - Spectral slope of the higher energy electron population
‘Host_Galaxy’ bytes17 - Best-fit host galaxy model
‘T_host’ float64 K Temperature of the host galaxy
‘chisq’ float64 - Chi square value of the fitting
Table 6: Description of table 1 (HDU 1).
Column Name Data Type Units Description
‘Source_Name’ bytes16 - Source name given in the 4FGL catalogue
‘nu_sync_peak’ float64 Hz Frequency of the synchrotron peak
‘nu_IC_peak’ float64 Hz Frequency of the inverse Compton peak
‘CD’ float64 - Compton dominance
‘jet_L_Sync’ float64 erg/s Jet luminosity due to the synchrotron component
‘jet_L_rad’ float64 erg/s Jet luminosity associated with radiative mechanisms
‘jet_L_B’ float64 erg/s Jet luminosity due to the magnetic field
‘jet_L_kin’ float64 erg/s Jet luminosity due to the electrons
‘jet_L_tot’ float64 erg/s Total jet luminosity
‘UB/Ue’ float64 - Ratio of magnetic to electron energy density
Table 7: Description of table 2 (HDU 2).
Column Name Data Type Units Description
‘Source_Name’ bytes16 - Source name given in the 4FGL catalogue
‘CTAO_significance’ float64 - CTAO expected detection significance
‘Redshift’ float64 - Redshift of the source
‘Redshift_reference’ bytes10 - Reference of the redshift estimate
Table 8: Description of table 3 (HDU 3).
Column Name Data Type Units Description
‘Source_Name’ bytes16 - Source name given in the 4FGL catalogue
‘log10nu’ float32 Hz log nu values used to reconstruct the broadband SED
‘log10nuFnu’ float32 erg/(cm2 s) log nuFnu values used to reconstruct the broadband SED
‘unc_log10nuFnu’ float32 erg/(cm2 s) Uncertainty in log10nuFnu
‘UL’ bool - Upper limits
‘instrument’ bytes37 - Instrument that took the corresponding SED data
Table 9: Description of table 4 (HDU 4).